We let T = the total change in the level of the lake. Then the problem translate to a sum. T= -5/8 + (-1/6)
Answer by stanbon(75887)

T= -5/8 + (-1/6)
---
T = [6*-5 + 8*-1]/(8*6)
----
T = -38/48
-----
T = -19/24
================
Cheers,
Stan H.
===============